OpenStudy (anonymous):

No one helps me! :O I give up on OpenStudy! what is the vertex of -16x^2+60x+16?

OpenStudy (anonymous):

using the formula \[x= \frac{ -b }{ 2a}\]

OpenStudy (anonymous):

@iambatman @confluxepic @TheSmartOne

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Can you please help me?

OpenStudy (preetha):

I am giving you some OwlBucks. Check your dm.

OpenStudy (mathstudent55):

\(y = ax^2 + bx + c\) \(y = -16x^2 + 60x + 16\) \(a = -16\); \(b = 60\); \(c = 16\) \(x = - \dfrac{b}{2a} \)